OpenStudy (syub):

Which points satisfy this linear quadratic system of equations? x+y=1 y=4x-3-x^2

OpenStudy (erak):

Set both equal to y, so you'd need to change the first one. x+y = 1 y = 1-x Now set both equations equal to each other and let it equal 0. 1-x = 4x-3-x^2 0 = -x^2 + 5x - 4 Put it in standard form, and now factor it.